OpenStudy (anonymous):

can anyone plz solve this???? fill in the blank with the numerator of the sun shown below x/x+1+x+5/x+1=??/x+1

OpenStudy (perl):

it looks like you are missing parentheses

OpenStudy (perl):

x / ( x + 1) + ( x + 5) / ( x + 1) = ? / (x+1) since we have a common denominator on the left side of the equation, you can add the numerators

OpenStudy (anonymous):

x / ( x + 1) + ( x + 5) / ( x + 1) = V / (x+1) ( x + x + 5) / ( x + 1) = V / (x+1) ( 2x + 5) / ( x + 1) = V / (x+1) 2x+5=V

OpenStudy (anonymous):

thank you guys :) !! it helped alot